# SciLitLLM: How to Adapt LLMs for Scientific Literature Understanding

### v3

@@ 4.1.1 Benchmarks
  • Instruction model benchmarks. We evaluate the instruct models on scientific literature understanding benchmarks: SciRIFF (Wadden et al., 2024) and SciAssess (Cai et al., 2024). Brief descriptions of them are provided in Appendix D.
@@ 4.1.2 Baselines
  • Base model baselines. We compare SciLitLLM-base against leading open-source base models: Llama-3.1 (AI@Meta, 2024), Qwen-2.5 (Team, 2024).
  • Instruction model baselines. We benchmark leading instruction LLMs including GPT-[3.5, 4o] (Brown et al., 2020; OpenAI, 2023), Llama3.1-[8B, 70B] (AI@Meta, 2024) and Qwen2.5-[7B, 14B, 72B] (Yang et al., 2024; Team, 2024) and Mistral-7B (Jiang et al., 2023). We also report the performance of SciTulu-7B (Wadden et al., 2024), which is a fine-tuned Llama2-7B (Touvron et al., 2023) model on SciRIFF.
@@ 4.2.1 Base model performance
The performance comparison of base models is shown in Table 2. SciLitLLM-base consistently outperforms other general base models across four scientific benchmarks. Specifically, compared with LLMs of less than 10 billion parameters, SciLitLLM-7B-Base shows an average accuracy improvement of 2.60% over Qwen2.5-7B. The results demonstrate that CPT on high-quality scientific corpora can significantly improve the LLMs’ ability to understand scientific contexts.
@@ 4.2.2 Instruct model performance
As shown in Table 3, SciLitLLM-7B-Instruct achieves the highest performance in all 4 domains on SciAssess, outperforming the second-best model (i.e., Llama3.1-8B) by 4.0%. Notably, on SciRIFF, it surpasses baseline models by a substantial margin of 10.1%. Additionally, SciLitLLM-14B-Instruct shows a 4.6% and 7.5% performance improvement over Qwen2-14B-Instruct on SciAssess and SciRIFF, respectively. It even outperforms other open-source models with five times more parameters (i.e., Llama3.1-70B and Qwen2.5-72B) on both benchmarks. This demonstrates the effectiveness of our proposed pipeline in adapting a general-purpose LLM to a specialized model for scientific literature understanding.
Detailed model performance on SciAssess is presented in Table 7, where SciLitLLM-7B lead in 15 out of 27 sub-tasks. Specifically, SciLitLLM excels in tasks such as table extraction and molecule generation, likely benefiting from the comprehensive task coverage in our synthetic instruction dataset – SciLitIns. On SciRIFF, SciLitLLM-7B and SciLitLLM-14B both rank first in 8 out of 11 evaluations [^8] In SciRIFF, the Qasper and SciFact tasks have two different evaluation methods and thus two results. For LLM-based evaluations in MuP and Qasper, we use GPT-4o instead of GPT-3.5, as was done in the original paper..
@@ 4.3.1 Scientific knowledge injection via CPT (Q2)
Model SciAssess SciRIFF
@@
Dataset SciAssess SciRIFF
@@
As shown in Table 6, applying SFT alone to the Qwen2.5-7B-Base model yields improvements on SciAssess (+3.3%) and SciRIFF (+6.7%). Further, incorporating CPT results in an extra 1.9% improvement on SciAssess and a 3.6% gain on SciRIFF. This shows the effectiveness of our CPT pipeline in curating high-quality CPT corpora and injecting scientific knowledge to LLMs.
@@ 4.3.2 SFT with SciLitIns (Q3)
We explore how fine-tuning with our synthetic instruction set, SciLitIns, impacts performance in scientific literature understanding. We incrementally add three datasets to the SFT training set: Infinity-Instruct, SciRIFF, and SciLitIns. As shown in Table 6, using only the general purpose instruction set (i.e., Infinity-Instruct) results in the lowest performance on both SciAssess (47.1%) and SciRIFF (51.2%). This indicates that fine-tuning LLMs on general instructions alone is insufficient for scientific literature understanding.
Adding the training set of SciRIFF to Infinity-Instruct improves performance on SciRIFF significantly (+5.5%) but decreases performance on SciAssess (-0.3%). This drop may be due to SciRIFF’s relatively narrower focus, primarily covering biomedicine and clinical medicine, while SciAssess spans broader fields such as biology, chemistry, and material science. This highlights the need for a diverse instruction set that covers diverse scientific domains. Finally, including SciLitIns boosts performance on both benchmarks, with SciAssess at 51.7% and SciRIFF at 60.6%. This demonstrates that SciLitIns, which is designed to contain instructions from less-represented scientific domains, is useful for improving model performance in scientific literature understanding.
@@ 4.3.3 Impact of Instruction filter
We conduct an ablation study to assess the impact of quality filter for synthetic instructions by varying whether the dataset SciLitIns  was filtered. As discussed in Section 3.2.2, this filter removes low-quality instructions evaluated from five key aspects. Table 6 shows that applying the filter improves the performance of SciLitLLM-7B on SciAssess (+0.8%) and SciRIFF (+0.5%). This demonstrates that SFT quality filtering process effectively excludes low-quality instructions, thereby boosting the performance of SciLitLLMon scientific literature understanding.
@@ 5 Limitations
  • Insufficient data volume. Compared with existing pre-training datasets (AI@Meta, 2024; Yang et al., 2024; Taylor et al., 2022), the amount of data used for CPT is comparatively small. Future work should consider incorporating a larger scientific corpus, potentially including scientific blogs or purely synthetic data.
@@ 6 Conclusion and Future Works
In this paper, we introduce SciLitLLM, a specialized model for scientific literature understanding. It is initialized with a general base model – Qwen2.5 (Yang et al., 2024), and trained through continual pre-training (CPT) and supervised fine-tuning (SFT). For effective scientific knowledge injection during CPT, we propose model-based format and grammar correction method, along with text quality filtering measures. To ensure high-quality and diverse instructions during SFT, we devise instruction synthesis and quality control approaches. Our experiments on widely-used benchmarks demonstrate the effectiveness of this pipeline in adapting a general model to the field of scientific literature understanding. Specifically, SciLitLLM-7B achieves a 4.0% improvement on the SciAssess (Cai et al., 2024) and a 10.1% improvement on the SciRIFF (Wadden et al., 2024) compared to leading models with fewer than 10 billion parameters. SciLitLLM-14B also surpasses leading open-source LLMs with around 70B parameters. We note that this pipeline could be easily adapted to other specialized domains, particularly those lacking adequate high-quality corpora and instruction sets.
@@
We have made extensive efforts to ensure the reproducibility of our work. Our proposed SciLitLLM models are available anonymously. The corresponding training pipeline have been described in Sections 3. The source code for all modules shown in Figure 5 is included in the supplementary material. The training datasets for both CPT and SFT stages are shown in Table 1. We will release the generated instruction sets SciLitIns after administrative procedure.
Appendix A Format & Grammar Correction Examples
@@
Appendix B CPT Quality Filter
@@
We train a Scientific Texts Quality Classifier on these labeled data samples. The classifier is a 109M BERT (Devlin et al., 2019) classifier, fine-tuned from the checkpoint of fineweb-edu-classifier (Anton et al., 2024). The model is trained for 20 epochs with a learning rate of 0.001 and a batch size of 1024. Ninety percent of the 50K samples are used as the training set, and the rest 10% are used as the validation set. The training process costs approximately 50 minutes on 4 A100 GPUs. We select the checkpoint from the epoch that yields the highest validation micro F1 score as our final checkpoint.
During Inference, we set batch size to 2048, and beam number to 1. The inference process costs 90 minutes on 4 A100 GPUs. We utilize the generated to filter out 25% data with the lowest quality. The distribution of the scores is demonstrated in Figure 6. The filtered-out 25% data are marked gray, while the remaining 75% CPT data are marked orange.
Figure 6: Score distribution of the CPT Data
Appendix C SFT Details
C.1 Instruction Generation Pipeline
@@ C.2 Instruction Deduplication
C.3 Quality Assessment of Generated SFT Instructions
@@
We show the quality statistics of synthetic instructions in Figure 7. Instructions with an average score of less than 4 are filtered out.
Figure 7: The quality of SciLitIns is evaluated from five aspects: clarity, complexity, correctness, usefulness, and adaptability (the higher the better). Instructions with an average score of less than 4 are filtered out.
Appendix D Benchmark Details
@@
Appendix E Detailed Performance on SciAssess
The detailed results on each tasks in SciAssess are shown in Table 7.

### v4

@@ 3.1.1 Benchmarks
  • Instruction model benchmarks. We evaluate the instruct models on scientific literature understanding benchmarks: SciRIFF (Wadden et al., 2024) and SciAssess (Cai et al., 2024). Due to the space limitation, we refer to their original paper for the detailed metrics of each task. Brief descriptions of them are provided in Appendix E.
@@ 3.1.2 Baselines
  • Base model baselines. We compare SciLitLLM-base against leading open-source base models: Llama-3.1 (AI@Meta, 2024), Qwen-2.5 (Qwen, 2024).
  • Instruction model baselines. We benchmark leading instruction LLMs including GPT-[3.5, 4o] (Brown et al., 2020; OpenAI, 2023), Llama3.1-[8B, 70B] (AI@Meta, 2024) and Qwen2.5-[7B, 14B, 72B] (Yang et al., 2024; Qwen, 2024) and Mistral-7B (Jiang et al., 2023). We also report the performance of SciTulu-7B (Wadden et al., 2024), which is a fine-tuned Llama2-7B (Touvron et al., 2023) model on SciRIFF.
@@ 3.2.1 Base model performance
The performance comparison of base models is shown in Table 2. SciLitLLM-base consistently outperforms other general base models across four scientific benchmarks, indicating its extensive domain knowledge. For instance, SciLitLLM-7B-Base achieves an average accuracy improvement of 2.60% over Qwen2.5-7B-Base, despite both models having a comparable parameter count. This improvement can be attributed to the specialized pretraining on high-quality domain-specific corpora. It benefits from exposure to scientific language patterns, terminology, and reasoning.
@@ 3.2.2 Instruct model performance
As shown in Table 3, compared with models with similar scale, SciLitLLM-7B achieves the highest performance in all 4 domains on SciAssess, outperforming the second-best model (i.e., Llama3.1-8B) by a notable 4.0%. On SciRIFF, it surpasses baseline models by an even more substantial margin of 10.1%. These results highlight the effectiveness of our pipeline in adapting a general-purpose LLM into a specialized model for scientific literature understanding.
Specifically, one key contributor is our synthetic instruction dataset — SciLitIns, which offers comprehensive task coverage across diverse domains. It ensures that the model can generalize across tasks with varying input data and reasoning complexities, such as table extraction and molecule generation. For example, SciLitLLM-7B leads in 15 out of 27 sub-tasks on SciAssess, indicating its strong ability to adapt to the specialized requirements of these tasks.
Notably, SciLitLLM-14B shows significant performance gains over Qwen2-14B on both benchmarks, with a 4.4% improvement on SciAssess and a 7.5% improvement on SciRIFF. These results suggest that our approach effectively leverages the larger capacity of the 14B parameter model, enabling it to encode more nuanced scientific reasoning and domain-specific knowledge. Interestingly, SciLitLLM-14B also outperforms much larger open-source models, such as Llama3.1-70B and Qwen2.5-72B, despite having only one-fifth the number of parameters. This demonstrates that our continual pretraining and instruction fine-tuning pipeline is not merely scale-dependent but strategically optimized for domain adaptation.
The performance on SciRIFF further illustrates this optimization. Comparing with open-source models with similar scale, both SciLitLLM-7B and SciLitLLM-14B rank first in 8 out of 11 evaluations, including challenging tasks like Qasper and SciFact. These tasks require advanced abilities in scientific question answering and fact verification, respectively, which are likely bolstered by the diversity of high-quality examples in SciLitIns.
These findings underscore the broader importance of designing domain-specific LLMs. While general-purpose models, such as Llama and Qwen, perform well on broader tasks, their lack of domain specialization limits their effectiveness in fields like scientific literature understanding. The results demonstrate that the adaptation pipeline designed by us can yield promising performance without relying solely on increased model size.
@@ 3.3.1 Scientific knowledge injection via CPT (Q2)
Benchmark SciAssess SciRIFF
Model
@@
Benchmark SciAssess SciRIFF
@@
As shown in Table 6, applying SFT alone to the Qwen2.5-7B-Base model yields improvements on SciAssess (+3.3%) and SciRIFF (+6.7%). This demonstrates that task-specific instruction fine-tuning effectively enhances performance by adapting the model to the specialized tasks in scientific literature understanding. Further, incorporating CPT results in an extra 1.9% improvement on SciAssess and a 3.6% gain on SciRIFF. These results highlight the unique role of CPT in pre-adapting the base model to scientific contexts. By curating a high-quality scientific corpus and using it for domain-specific pretraining, the CPT stage injects foundational scientific knowledge into the model. This foundational understanding reduces the gap between general-purpose pretraining and the specific requirements of downstream tasks, allowing the subsequent SFT stage to focus on fine-grained task adaptation rather than compensating for domain knowledge deficiencies.
@@ 3.3.2 SFT with SciLitIns (Q3)
We explore how fine-tuning with our synthetic instruction set, SciLitIns, impacts performance in scientific literature understanding. We incrementally add three datasets to the SFT training set: Infinity-Instruct, SciRIFF, and SciLitIns. As shown in Table 6, using only the general purpose instruction set, Infinity-Instruct, results in the lowest performance on both SciAssess (47.1%) and SciRIFF (51.2%). This result highlights the limitations of general-purpose instructions in addressing the nuanced demands of scientific literature tasks. General-purpose datasets often lack the domain-specific structure, reasoning, and terminology required for tasks like scientific fact verification or molecule generation. These findings emphasize the need for specialized instructions to bridge the gap between general capabilities and domain-specific expertise.
Adding the training set of SciRIFF to Infinity-Instruct improves performance on SciRIFF significantly (+5.5%) but decreases performance on SciAssess (-0.3%). This drop may be due to SciRIFF’s relatively narrower focus, primarily covering biomedicine and clinical medicine, while SciAssess spans broader fields such as biology, chemistry, and material science. This highlights the need for a diverse instruction set that covers diverse scientific domains. Replacing SciRIFF with SciLitIns in the fine-tuning process yields better performance than Infinity-Instruct on both benchmarks, achieving 50.2% on SciAssess and 54.8% on SciRIFF. The diverse task coverage in SciLitIns plays a key role in this improvement. By incorporating synthetic instructions from underrepresented domains, SciLitIns broadens the model’s exposure to scientific knowledge, enabling it to generalize across varied scientific disciplines.
Finally, combining SciRIFF and SciLitIns yields the best results, with SciAssess reaching 51.7% and SciRIFF achieving 60.6%. These gains suggest that SciLitIns and SciRIFF complement each other, addressing the limitations of each dataset individually. While SciRIFF focuses on highly specific domains like biomedicine, SciLitIns fills the gaps by covering tasks from broader and less-represented scientific areas.
@@ 3.3.3 Impact of Instruction filter
We conduct an ablation study to assess the impact of quality filter for synthetic instructions by varying whether the dataset SciLitIns  was filtered. As discussed in Section 2.2.2, this filter removes low-quality instructions evaluated from five key aspects. Table 6 shows that applying the filter improves the performance of SciLitLLM-7B on SciAssess (+0.8%) and SciRIFF (+0.5%).
While the performance improvements from filtering are evident, the gains are relatively modest. This suggests that the dataset was already of reasonably high quality before filtering, with only a small proportion of low-quality instructions being removed. It also indicates that the fine-tuning process is robust enough to tolerate some noise, although further improvements can be achieved with higher-quality inputs.
@@ 4 Limitations
  • Insufficient data volume. Compared with existing pre-training datasets (AI@Meta, 2024; Yang et al., 2024; Taylor et al., 2022), the amount of data used for CPT is comparatively small. While the dataset has been carefully curated to ensure quality and domain relevance, the limited size poses challenges in achieving comprehensive representation across diverse scientific disciplines. With a smaller corpus, certain specialized may be underrepresented, limiting the model’s ability to perform well on tasks within these areas. While increasing dataset size is a logical step, it is crucial to maintain high-quality filter to avoid introducing noise. The effectiveness of the current CPT stage demonstrates that a smaller but well-curated dataset can still deliver significant gains. Therefore, the challenge lies in scaling the corpus size while preserving its relevance and quality.
@@ 5 Conclusion and Future Works
In this paper, we introduce SciLitLLM, a specialized model for scientific literature understanding. It is initialized with a general base model – Qwen2.5, and trained through continual pre-training (CPT) and supervised fine-tuning (SFT). For effective scientific knowledge injection during CPT, we propose model-based format and grammar correction method, along with text quality filtering measures. To ensure high-quality and diverse instructions during SFT, we devise instruction synthesis and quality control approaches. Our experiments on widely-used benchmarks demonstrate the effectiveness of this pipeline in adapting a general model to the field of scientific literature understanding. Specifically, SciLitLLM-7B achieves a 4.0% improvement on the SciAssess (Cai et al., 2024) and a 10.1% improvement on the SciRIFF (Wadden et al., 2024) compared to leading models with fewer than 10 billion parameters. SciLitLLM-14B also surpasses leading open-source LLMs with around 70B parameters. We note that this pipeline could be easily adapted to other specialized domains, particularly those lacking adequate high-quality corpora and instruction sets.
@@
We have made extensive efforts to ensure the reproducibility of our work. Our proposed SciLitLLM models are available. The corresponding training pipeline have been described in Sections 2. The source code for all modules shown in Figure 5 is included in the supplementary material. The training datasets for both CPT and SFT stages are shown in Table 1.
Acknowledgement
The work of Sihang Li, Yaorui Shi, and Xiang Wang, is in part supported by the National Natural Science Foundation of China (92270114).
Appendix A Related Works
A.1 Knowledge Injection via Continual Pre-training
Pre-training a language model is usually conducted on a large corpus of textual data to learn the statistical properties of language (Radford et al., 2019; Brown et al., 2020; Lo et al., 2020; Soldaini et al., 2024). To further inject domain knowledge into a general LLM after pre-training, researchers engage in continual pre-training (CPT) on high-quality domain-specific corpora (Jin et al., 2022; Sun et al., 2020), sometimes combined with general corpora. This process enhances the model’s fundamental understanding abilities in specific downstream domains while mitigating catastrophic forgetting of general knowledge (Wu et al., 2022; Ke et al., 2023; Mehta et al., 2023). See the comprehensive study (Gupta et al., 2023) for different warm-up strategies for CPT. Additionally, the CPT corpora can be augmented by transforming them into an instruction-response format (Cheng et al., 2023; 2024). Furthermore, the scaling law (Hoffmann et al., 2022) of domain-specific CPT (Que et al., 2024) is explored to determine the optimal mix of data. However, these studies primarily focus on training dynamics and data recipes, leaving the pre-processing for scientific data, especially raw PDF files, largely unexplored. Exhibiting such steps is essential for generating high-quality domain corpora and effectively injecting domain knowledge, representing a significant challenge for practitioners.
A.2 Domain Adaptation via Supervised Fine-tuning
Supervised fine-tuning (SFT) modifies a pre-trained language model to follow specific instructions or perform designated tasks by fine-tuning it on a targeted, task-specific dataset (Raffel et al., 2020; Wei et al., 2021; Liu et al., 2023; Li et al., 2024; Liu et al., 2024; Liao et al., 2024; 2025). Applying SFT to a general LLM for specific domain adaptation has demonstrated effectiveness in various fields: in medicine (Clusmann et al., 2023), corpora of medical literature and clinical notes are used; in law (Cui et al., 2024), legal documents and case law are compiled; and in finance (Wu et al., 2023), financial reports and market data are utilized. In the scientific domain, several studies have specialized LLMs for scientific tasks, often necessitating the construction of a substantial domain-specific dataset with SFT. For example, SciGLM (Zhang et al., 2024a) leverages existing LLMs to generate step-by-step reasoning for unlabelled scientific instructions. ChemLLM (Zhang et al., 2024b), a more specified LLM in the chemistry field, collects structured chemical data from a vast selection of online databases and transforms this structured data into a question-answering format for SFT. SciRIFF (Wadden et al., 2024) converts existing literature understanding datasets into natural language input-output pairs suitable for instruction-tuning using pre-defined templates. However, benchmark studies (Feng et al., 2024; Cai et al., 2024) indicate that SFT alone may not provide adequate scientific knowledge to excel in relevant tasks. This suggests the need for a more comprehensive approach that combines domain knowledge infusion with instruction-following enhancements.
A.3 LLMs for Scientific Literature Understanding
In the scientific domain (Luo et al., 2024; Liu et al., 2025; Shi et al., 2024), existing strategies for developing specialized LLMs mostly fall into two categories: (1) Supervised fine-tuning with scientific instructions. This approach requires a large, high-quality, and diverse set of instructions to cultivate problem-solving abilities for scientific tasks. Representative works (e.g., SciGLM (Zhang et al., 2024a), ChemLLM (Zhang et al., 2024b), and SciRIFF (Wadden et al., 2024)) have been detailed in Section A.2. (2) Pre-training with scientific corpora. This approach involves pre-training on a large corpus of scientific texts to improve performance on downstream scientific tasks. Early attempts, such as SciBert (Beltagy et al., 2019) and KV-PLV (Zeng et al., 2022), are based on BERT (Devlin et al., 2019) and pre-trained on a large corpus of scientific text for downstream scientific task enhancement. More recently, Galactica (Taylor et al., 2022) is pre-trained on a vast corpus of scientific literature, including research papers, scientific articles, and other relevant scientific texts. Despite these advances, two major limitations hinder these models from excelling in scientific literature understanding: (1) lack of scientific knowledge, and (2) inability to follow task instructions. To address these challenges, we propose a combined pipeline of CPT and SFT to devise a specialized LLM for scientific literature understanding. It injects domain-specific knowledge through CPT while enhancing task-specific instruction-following abilities through SFT, leading to a more capable LLM for scientific literature understanding.
Appendix B Format & Grammar Correction Examples
@@
Appendix C CPT Quality Filter
@@
We train a Scientific Texts Quality Classifier on these labeled data samples. The classifier is a 109M BERT (Devlin et al., 2019) classifier, fine-tuned from the checkpoint of fineweb-edu-classifier (Anton et al., 2024). The model is trained for 20 epochs with a learning rate of 0.001 and a batch size of 1024. Ninety percent of the 50K samples are used as the training set, and the rest 10% are used as the validation set. The training process costs approximately 50 minutes on 4 A100 GPUs. We select the checkpoint from the epoch that yields the highest validation micro F1 score as our final checkpoint. During Inference, we set batch size to 2048, and beam number to 1. The inference process costs 90 minutes on 4 A100 GPUs. We utilize the generated to filter out 25% data with the lowest quality.
We also conduct empirical experiments to investigate the sensitivity of model performance to the quality filtering threshold. The results in Table 7 show that setting the filtering threshold at 25% achieves the best overall performance, effectively injecting scientific domain knowledge into the model. When the threshold is set too high (e.g., 35%), the average quality of the corpus improves; however, the significant reduction in data volume limits the ability to fully capture domain knowledge. Conversely, when the threshold is set too low (e.g., 15%), the increase in data volume includes more low-quality content, which negatively impacts model performance. Thus, the trade-off between data quality and quantity is crucial.
Filtering Threshold MMLU-Pro-Bio MMLU-Pro-Chem MMLU-Pro-Heal MaScQA Avg.
15% 70.59 50.32 53.14 53.14 58.32
25% 70.45 51.21 54.06 54.06 58.75
35% 69.82 49.52 51.64 61.64 57.67
Table 7: 7B base model performance (accuracy %) with different CPT data filtering thresholds.
Appendix D SFT Details
D.1 Instruction Generation Pipeline
@@ D.2 Instruction Deduplication
D.3 Quality Assessment of Generated SFT Instructions
@@
Appendix E Benchmark Details
@@
Appendix F Detailed Performance on SciAssess
The detailed results on each task in SciAssess are shown in Table 8.
@@ Appendix F Detailed Performance on SciAssess
Appendix G Contamination Study
Eval Datasets In-house Textbooks (%) In-house Journals (%)
SciRIFF 1.1 0.7
SciAssess 11.0 1.1
Table 9: Contamination rates in textbook and journal datasets.
Category Contamination Rates (%) SciLitLLM-7B SciLitLLM-14B
Full Set (%) Clean Set (%) Full Set (%) Clean Set (%)
Biology 12.2 65.9 66.3 67.5 67.6
Chemistry 13.3 54.4 55.8 64.2 65.2
Material 12.5 52.5 50.7 61.8 62.9
Medicine 8.7 33.6 34.2 42.4 43.9
Mean 12.0 51.6 51.6 59.0 59.8
Table 10: Performance comparison of SciLitLLM on full and clean evaluation sets of SciAssess.
To investigate data contamination in our datasets and its potential impact on downstream performance, we conducted a contamination analysis in this section. Prior studies consistently indicate that data contamination in pre-training datasets has negligible effects on evaluation performance:
  • The GPT-4 technical report states: Contamination overall has very little effect on the reported results (OpenAI, 2023). For instance, tasks such as AP US History (73% contamination rate), AP World History (47% contamination rate), and LSAT (39% contamination rate) show negligible differences in performance between the full evaluation set and a clean evaluation set with all contaminated data points removed.
  • PaLM analyzed datasets with contamination rates ranging from 20% to 75% and concluded that data contamination does not cause meaningful inflation of our reported results (Chowdhery et al., 2023).
For our contamination analysis, we adopt the contamination detection method described in the GPT-4 technical report: For each evaluation instance, we randomly extract three substrings of 50 characters each. A match is determined if any of the three sampled substrings appears as a part of the processed training example.
The contamination rates for our textbook and journal datasets are summarized in Table 9. Our analysis reveals that contamination rates are very low for SciRIFF but higher for SciAssess, particularly in the textbook dataset. To evaluate whether contamination impacts model performance, we assessed SciLitLLM on SciAssess using both the full evaluation set and a clean evaluation set (with all contaminated examples removed). As shown in Table 10, the differences in performance between the full and clean evaluation sets are minimal, consistent with prior findings. These results reinforce the conclusion that data contamination does not significantly affect the performance of SciLitLLM.
Appendix H Human Evaluation of Dataset Quality
To further validate the effectiveness of the proposed framework for creating the CPT and SFT datasets, we conducted a human evaluation study. This study involved sampling random subsets of unfiltered CPT and SFT data and comparing human annotations to the scores generated by our quality filters.
For the CPT dataset, we sampled 50 entries from the textbook and journal datasets. For the SFT dataset, we sampled 50 entries from SciLitIns. Each sampled entry was independently evaluated by four annotators using the same prompts provided to the LLMs. To assess the alignment between human annotations and the quality filter, we computed two measures:
  • Human-Human Agreement: Calculated as the average Spearman correlation coefficient across all pairs of annotators.
  • Human-Quality Filter Agreement: Calculated as the average Spearman correlation between each annotator’s scores and the quality filter’s score.
A higher Spearman correlation indicates a stronger agreement between the compared sets of scores. The evaluation results are summarized in Table 11. The results show that for both datasets, the Human-Quality Filter Agreement is comparable to, or even exceeds, the Human-Human Agreement. This indicates that the quality filter closely aligns with human annotations, demonstrating its reliability in capturing quality as perceived by humans.
Agreement Human-Human Human-Quality Filter
CPT 0.58 0.76
SFT 0.89 0.88
Table 11: Comparison of Human-Human Agreement and Human-Quality Filter Agreement for CPT and SFT datasets.